OpenStudy (ebonyserrano):

Three consecutive integers have a sum of 87. Find the integers.

OpenStudy (3mar):

Any ideas?

OpenStudy (steve816):

Heyyy let's do some fun math! If three consecutive integers have a sum of 87, it would look like\[\huge x + (x + 1) + (x + 2) = 87\] Do you think you can solve for x?

OpenStudy (taliam15):

I just sort of started somewhere and worked from there. I didn't even think in terms of a formula :/ Pick three numbers that are consecutive and add them together.

OpenStudy (ebonyserrano):

28

OpenStudy (ebonyserrano):

thats what i got for x

OpenStudy (taliam15):

Perfect, now plug in what you got for x into the binomials and you'll get the integers :)

OpenStudy (taliam15):

Does that make sense? @steve816 did a great job with that equation

OpenStudy (steve816):

That is correct! Now as @taliam15 said, plug in 28 for x to get the 3 values!

OpenStudy (steve816):

Welcome to OpenStudy @taliam15 !!!

OpenStudy (taliam15):

Thank you :)

OpenStudy (taliam15):

@Ebonyserrano did we help at all?

OpenStudy (ebonyserrano):

yes

OpenStudy (ebonyserrano):

thank you guys

OpenStudy (taliam15):

okay good :D and no problem glad I could finally assist someone in some way on here lol

OpenStudy (ebonyserrano):

i just got 87

OpenStudy (ebonyserrano):

so is it 28, 87, and 87?

OpenStudy (steve816):

Actually, it would be 28, 29, and 30.

OpenStudy (taliam15):

oh, no. Okay so you're going to plug in 28 for each value of x, even the ones in the parenthesis.

OpenStudy (ebonyserrano):

i did

OpenStudy (taliam15):

yeah, and when you do that you get what he said ^

OpenStudy (ebonyserrano):

ok thanks

OpenStudy (taliam15):

You sure you got it?

OpenStudy (ebonyserrano):

can you guys help me with two more?

OpenStudy (taliam15):

I can show you

OpenStudy (taliam15):

\[\left( 28 \right)\left( 28 + 1 \right)\left( 28+2) \right)\]

OpenStudy (taliam15):

28 + 29 + 30 = 87

OpenStudy (taliam15):

oh darn I have somewhere to go for about an hour if you wanna hold out I'll help but maybe @steve816 could help, sorry

OpenStudy (ebonyserrano):

ok thanks